In 4 days your family drive 5/7 of a trip. Your rate of travel is the same throughout the trip. The total trip is 1250 miles. How many more days until you reach your destination?
Answer by Nfrey78(18)

In 4 days your family drive 5/7 of a trip. Your rate of travel is the same throughout the trip. The total trip is 1250 miles. How many more days until you reach your destination?
Find the amount traveled in one day by dividing
(5/7)/4 = 5/28 of the trip
How many days would it take to add up to 1
1/(5/28) = 5.6 days total
since we already traveled for 4 days, we have 1.6 more days to go